Lower-of-Cost-Net-Realizable-Value Method The following data are taken from the Browning Corporations inventory accounts: ItInventory Turnover and Days Sales in Inventory The Western Company installed a new inventory management system at the beginnThe LIFO Inventory Reserve Southeast Steel Company uses the LIFO inventory costing method to value its ending inventory. TheInventory Costing Methods—Perpetual Method The Luann Company uses the perpetual inventory system. The following July data are
Answer #1

1.

ACE = 100 * (Lower of 27 or 25) = 2500

BDF = 300 * (Lower of 29 or 31) = 8700

GHJ = 400 * (Lower of 22 or 18) = 7200

MBS = 200 * (Lower of 23 or 27) = 4600

Ending inventory value = 2500+8700+7200+4600

= 23,000